# references.bib

### v2

@@ 4.1 Models
Lang et al. ($13$M parameters). The architecture developed in [Lang2023] for CH estimation, but trained (from scratch) for AGB estimation, with standard $\ell_{2}$-loss (rather than the original log-likelihood loss). It consists of a series of residual blocks with separable convolutions without any downsampling.
@@ 4.2 Results
CH RGBN S2 (all) LC DEM P2 GBDT FCN UNet Lang et al. [Lang2023]
@@ 5 Discussion
Results. Our various analyses indicate that including input features beyond Sentinel-2 is beneficial for the task of AGB estimation, and that models with more parameters tend to better capture the complex relationship between input features and AGB.
@@ 8.1 Dataset description
Each sample in the dataset contains a pair of pre-cropped, pre-normalized images along with their corresponding AGB labels.
The load_dataset function provides the following configuration options:
  • normalize_data: {True, False}
    Whether to return normalized (0-1) data or raw values.
  • additional_features: {}
    A list of additional features the dataset should include. Refer to the documentation below for more details.
  • patch_size: 25
    The size of the returned patch (in pixels). The maximum value is 25 pixels, which corresponds to 250 meters.
Each image consists of 24 channels, organized into the following categories:
  • Spectral Bands:
    B01, B02, B03, B04, B05, B06, B07, B08, B8A, B09, B11, B12
  • Geographical Coordinates:
    lat_cos, lat_sin, lon_cos, lon_sin, where:
@@
  • ALOS PALSAR Bands:
    alos_hh, alos_hv
  • Canopy Heights:
    ch, ch_std
  • Land Cover Information:
    lc_cos, lc_sin, lc_prob, where:
@@
    lc_prob $=lc\_prob/100$
  • Digital Elevation Model (DEM):
    dem
The channels are ordered as follows: (Spectral Bands) | (Geographical Coordinates) | (ALOS PALSAR Bands) | (Canopy Heights) | (Land Cover Information) | DEM. In more details, this corresponds to the following bands: (B01 B02 B03 B04 B05 B06 B07 B08 B8A B09 B11 B12) | (lat_cos, lat_sin, lon_cos, lon_sin) | (alos_hh, alos_hv) | (ch, ch_std) | (lc_cos, lc_sin, lc_prob) | dem.
Users can include a list of additional features from the options below in their dataset configuration:
  • "agbd_se" - AGBD Standard Error: The uncertainty estimate associated with the aboveground biomass density prediction for each GEDI footprint.
  • "elev_lowes" - Elevation: The height above sea level at the location of the GEDI footprint.
  • "leaf_off_f" - Leaf-Off Flag: Indicates whether the measurement was taken during the leaf-off season, which can impact canopy structure data.
  • "pft_class" - Plant Functional Type (PFT) Class: Categorization of the vegetation type (e.g., deciduous broadleaf, evergreen needleleaf).
  • "region_cla" - Region Class: The geographical area where the footprint is located (e.g., North America, South Asia).
  • "rh98" - RH98 (Relative Height at 98%): The height at which 98% of the returned laser energy is reflected, a key measure of canopy height.
  • "sensitivity" - Sensitivity: The proportion of laser pulse energy reflected back to the sensor, providing insight into vegetation density and structure.
  • "solar_elev" - Solar Elevation: The angle of the sun above the horizon at the time of measurement, which can affect data quality.
  • "urban_prop" - Urban Proportion: The percentage of the footprint area that is urbanized, helping to filter or adjust biomass estimates in mixed landscapes.
  • "gedi_num_days" - Date of GEDI Footprints: The specific date on which each GEDI footprint was captured, adding temporal context to the measurements.
  • "s2_num_days" - Date of Sentinel-2 Image: The specific date on which each Sentinel-2 image was captured, ensuring temporal alignment with GEDI data.
  • "lat" - Latitude: Latitude of the central pixel.
  • "lon" - Longitude: Longitude of the central pixel.
@@ 8.2 Hosting, licensing, and maintenance plan
The dataset will be maintained by the authors to the best of their abilities and availability. The authors bear all responsibility in case of violation of rights.

### v3

@@ 4.1 Models
Lang et al. ($13$M parameters). The architecture developed in [Lang2023] for CH estimation, but trained (from scratch) for AGB estimation, with standard $\ell_{2}$-loss (rather than the original log-likelihood loss). It consists of a series of residual blocks with separable convolutions without any downsampling. We use $728$ filters per block (instead of $256$).
@@ 4.2 Results
CH RGBN S2 (all) LC DEM P2 GBDT FCN UNet Lang et al.
@@ 5 Discussion
Results. Our various analyses indicate that including input features beyond Sentinel-2 is beneficial for the task of AGB estimation, and that more complex models tend to better capture the complex relationship between input features and AGB.
@@ 8.1 Dataset description
Each sample in the dataset contains a pair of pre-cropped images (patches) and their corresponding AGB labels. You can find an up-to-date description of the dataset on HuggingFace.
We hereby provide more information on the encoding of certain features:
  • Geographical Coordinates — we perform sine/cosine encoding, as follows:
@@
  • Land Cover Information — we perform sine/cosine encoding, as follows:
@@
    where lc_class is the discrete classification of the land cover.
Additionally from the features described in this manuscript, users of the dataset have access to the following features (mainly GEDI L4A metadata):
  • AGBD Standard Error: The uncertainty estimate associated with the aboveground biomass density prediction for each GEDI footprint.
  • Elevation: The height above sea level at the location of the GEDI footprint.
  • Leaf-Off Flag: Indicates whether the measurement was taken during the leaf-off season, which can impact canopy structure data.
  • Plant Functional Type (PFT) Class: Categorization of the vegetation type by GEDI (e.g., deciduous broadleaf, evergreen needleleaf).
  • Region Class: The geographical area where the footprint is located (e.g., North America, South Asia).
  • RH98 (Relative Height at 98%): The height at which 98% of the returned laser energy is reflected, a key measure of canopy height.
  • Sensitivity: The proportion of laser pulse energy reflected back to the sensor, providing insight into vegetation density and structure.
  • Solar Elevation: The angle of the sun above the horizon at the time of measurement, which can affect data quality.
  • Urban Proportion: The percentage of the footprint area that is urbanized, helping to filter or adjust biomass estimates in mixed landscapes.
  • Date of GEDI Footprints: The specific date on which each GEDI footprint was captured, adding temporal context to the measurements.
  • Date of Sentinel-2 Image: The specific date on which each Sentinel-2 image was captured, ensuring temporal alignment with GEDI data.
@@ 8.2 Hosting, licensing, and maintenance plan
The dataset will be maintained by the authors to the best of their abilities and availability.
@@ 8.3 GBDT model initialization parameters
8.4 Dense predictions
We generate the dense $10$m resolution AGB estimates for all Sentinel-2 tiles in the dataset coverage. For each tile, we pick the Sentinel-2 product that has the least cloud coverage over all products available, and perform inference on that product. As such, our dense predictions span multiple years. On the ETH Research Collection, the file Inference > mapping.pkl indicates which specific product was used for inference for each tile. Note that this was a design choice, and one can generate dense predictions for specific years.